# MLPerf Power: Benchmarking the Energy Efficiency of Machine Learning Systems from µWatts to MWatts for Sustainable AI

### v1

@@ Abstract
Rapid adoption of machine learning (ML) technologies has led to a surge in power consumption across diverse systems, from tiny IoT devices to massive datacenter clusters. Benchmarking the energy efficiency of these systems is crucial for optimization, but presents novel challenges due to the variety of hardware platforms, workload characteristics, and system-level interactions. This paper introduces MLPerf® Power, a comprehensive benchmarking methodology with capabilities to evaluate the energy efficiency of ML systems at power levels ranging from microwatts to megawatts. Developed by a consortium of industry professionals from more than 20 organizations, MLPerf Power establishes rules and best practices to ensure comparability across diverse architectures. We use representative workloads from the MLPerf benchmark suite to collect 1,841 reproducible measurements from 60 systems across the entire range of ML deployment scales. Our analysis reveals trade-offs between performance, complexity, and energy efficiency across this wide range of systems, providing actionable insights for designing optimized ML solutions from the smallest edge devices to the largest cloud infrastructures. This work emphasizes the importance of energy efficiency as a key metric in the evaluation and comparison of the ML system, laying the foundation for future research in this critical area. We discuss the implications for developing sustainable AI solutions and standardizing energy efficiency benchmarking for ML systems.
@@ I Introduction
The advancement in ML capabilities has coincided with a notable increase in concerns about the power consumption of ML systems [69, 11, 68, 71]. As the ecological impact and operational expenses of these systems continue to increase, they have become a significant concern for both industry and academia [75, 66, 73, 15]. The critical nature of this challenge highlights the need for a standardized and thorough benchmarking approach to measure the energy efficiency of ML systems, from small IoT devices to large-scale cloud infrastructures. This method must also guarantee accurate reproducibility, facilitate meaningful comparisons, and encourage advances in energy-efficient ML technologies.
@@
These diverse requirements emphasize the need for a comprehensive power measurement methodology that is inherently adaptable. To address the multifaceted challenges of measuring and comparing the energy efficiency of ML systems, we developed MLPerf Power, a comprehensive benchmarking methodology. This initiative is the result of collaborative efforts among 20 organizations actively participating in the MLPerf Training and Inference [64] benchmarks, which have long served as the industry standard for measuring ML system performance. MLPerf Power extends this framework, recognizing that performance metrics alone are insufficient to enable the development of sustainable AI systems for the future.
@@ II Background, Motivation and Prior Work
This section establishes the requirements for benchmarking and measuring the energy efficiency of ML systems and demonstrates how our approach advances the field. We outline the key requirements and then gives an overview of the limitations of existing methods using Table I. MLPerf Power was specifically designed to meet all the identified requirements
@@ III ML System Power Measurement Challenges
This section outlines the challenges of creating a standardized approach to evaluate power consumption in ML systems.
@@ III-A Scale of ML Systems
Figure 2: The power consumption range across MLPerf categories, highlighting the need for scalable power measurement.
@@
MLPerf supports a large diversity of submitters, each with their own telemetry and monitoring setup. However, measurement of power between connections is challenging, as switches may not have as much telemetry to report power as compute servers, and cloud service providers often do not want to reveal their network topology to competitors. Measuring cooling power fairly is another challenge. Air-cooled servers have fans whose power is measured when measuring node power, but direct liquid cooled or immersion cooled systems pay their cooling cost at the datacenter level, making it difficult to attribute what fraction was used by the ML job under test. To address these challenges, we need to develop a software-based approach that leverages existing telemetry systems, allowing for comprehensive power measurement without compromising security or requiring extensive hardware modifications.
@@
Mobile MLPerf Mobile benchmarks the performance of mobile ML systems [30]. However, we do currently support mobile systems power measurement in MLPerf Power due to the complexity and variability of smartphones. Mobile devices comprise a wide range of components, including communication interfaces, sensors, and peripherals, each with significant fluctiations in power contumption patterns and active components[56]. Additionally, the battery-powered nature of these devices means they cannot be simply connected to an external power source for measurement without altering normal operation. Thermal throttling further complicates the relationship between power consumption and performance metrics [61]. Given these complexities, establishing a fair and transparent benchmark for measuring power consumption in battery-powered devices is beyond the current scope of work.
@@ IV MLPerf Power Measurement Methodology
This section outlines the core concepts, metrics, and philosophy behind the MLPerf power measurement approach.
@@ IV-A Common Principles Across All Segments
First, a fundamental tenet of MLPerf Power is the emphasis on measuring full system power consumption. This approach recognizes that ML workloads involve various components beyond the primary computing units. In datacenter systems, we account for compute nodes, interconnect fabric, and associated cooling infrastructure. For edge devices, we consider the entire system-on-chip, including any peripherals that cannot be powered down during operation. In tiny systems, we measure the power of the entire device, including any always-on components. Figure 3 illustrates our measurement for the different scales of systems. Our measurements closely reflect the true energy cost of running ML workloads in real-world scenarios, rather than focusing solely on idealized component-level efficiencies.
Second, ML workloads typically consist of multiple phases: initialization, execution, and result assimilation. The MLPerf Power methodology emphasizes the precision of aligning power measurements with these phases, particularly focusing on the execution phase. This alignment is crucial because it allows for accurate attribution of power consumption to specific parts of the workload, ensures comparability across different systems, and provides insight into the energy efficiency of the core ML computations, separate from setup or teardown operations. To achieve this in tiny systems, we use hardware pins to signal the start and end of inference. In larger systems, we use software-based timestamps to demarcate the execution phase. For training workloads, we parse performance logs to determine the start and stop times of the run.
Third, unlike evaluating overall system performance, the execution phase is more important than end-to-end runtime in power measurement considerations. The execution phase represents the core ML computations, which are typically the most energy-intensive part of the workload. It also provides the most relevant power measurement data to draw actionable insights to improve the energy efficiency of ML systems where it matters most. For inference tasks we measure the power during the actual inference computations. For training tasks, we focus on the power consumed during the training iterations, excluding the setup or data preparation steps. This approach, combined with full system power measurement and precise alignment with workload phases, forms the foundation of MLPerf Power’s methodology across all segments, from the smallest IoT devices to the largest datacenter installations.
@@
Finally, we employ a unified approach to evaluating energy efficiency on different scales and types of systems. For throughput benchmarks, such as datacenter and offline edge inference, we measure throughput in Samples/s and power consumption in Watts. We calculate energy efficiency as (Samples/s)/(Watts), yielding Samples/Joule. For latency benchmarks such as tiny inference [5], we use the inverse of energy in 1/Joules. We treat these two metrics as comparable, despite their different units, because the benchmarks fix different execution factors and thus require different performance evaluations. This standardized approach allows for meaningful comparisons across the diverse range of systems.
@@ IV-B MLPerf Inference Setup and Operation
To accurately measure such low-power devices, we utilize an I/O Manager that sits between the host and the System Under Test (SUT). This setup prevents parasitic powering, where the SUT might inadvertently draw power from the host’s signal lines, skewing the measurements. The I/O Manager provides a USB connection to the host and connects to the SUT’s UART through an isolating level shifter, ensuring minimal power transfer. We capture detailed energy waveforms to analyze power consumption patterns. The SUT toggles a pin to indicate the start and end of the inference, allowing for a precise measurement of the execution phase. To mitigate the impact of slow interfaces, such as UART, on measurements, we repeat the inference multiple times and average the results.
@@
Currently, the SUT collects performance log files that timestamp different phases of execution. To help us improve measurement accuracy, we employ a range mode that involves an initial run to determine the maximum current and voltage levels for a particular workload. Subsequent runs then use fixed analyzer ranges based on these observed peaks, allowing for higher accuracy within that range. For multi-node inference configurations, we adapt our methodology to handle multiple analyzers connected to a single SUT or multiple SUTs connected to a single analyzer. This flexibility allows us to accurately measure power consumption in more complex distributed inference scenarios.
@@ IV-C MLPerf Training and HPC Setup and Operation
Power measurement at the node level is the foundation of our approach. For each participating compute node, power measurements are taken using the submitter’s own telemetry systems, such as IPMI [46] or RedFish [20]. We recommend out-of-band measurement techniques to minimize interference with the ML job itself. In cases where individual node measurement is not feasible, power can be measured at the Power Distribution Unit (PDU) level, provided that all nodes supplied by that PDU are involved in the ML job. Submitters are required to thoroughly document the accuracy of their telemetry systems for review.
@@ IV-D Validation and Reporting
Quality Accuracy requirements form the foundation of our validation process. Although we recognize that the diverse range of systems and scales in ML workloads requires different approaches to power measurement, all must meet stringent accuracy standards as described by MLPerf Inference and Training [64, 44]. For datacenter and high-performance computing systems, we require documentation of the telemetry systems used, including their accuracy specifications and calibration procedures. Edge systems using external power analyzers must employ SPEC-approved devices, known for their high accuracy and reliability. For TinyML, we mandate specialized micro-power instruments with target accuracy ratings [5].
@@
The HPC and training reporting and submission process is designed to more comprehensive, flexible, and contextually rich. Submitters are required to report power consumption data along with additional information that could affect these measurements, including details on cooling solutions, power management techniques, and environmental conditions during the testing. The process allows for reporting at either the PDU or node level, accommodating different measurement setups. Submitters run the benchmark with their chosen telemetry option, estimate switch power if necessary, and post-process power logs to conform to the MLPerf Logging Library format. For systems where certain components’ power consumption must be estimated, such as interconnect power in large-scale training sets, dedicated sections are provided in the reporting format to document the estimation methodology.
Review Our review process is rigorous and multilayered, designed to catch any inconsistencies or anomalies in the submissions. Upon receiving a submission, our review committee, composed of experts in ML systems and power measurement, examines the reported data and documentation. They verify that all required information is present, that the reported measurements comply with the accuracy requirements, and that the power consumption figures are consistent with the system specifications and workload characteristics.
@@ V Results/Evaluation
We analyze this diverse range of data to gain an understanding of comparative trends across different workloads and system scales, shedding light on the areas of focus within the industry. MLPerf workloads span a diverse range of ML tasks, from resource-constrained keyword spotting to complex computer vision, sophisticated recommendation systems, and state-of-the-art large language models (LLMs) [45, 47].
@@ V-C Workload-Specific Insights
The energy footprint of each benchmark provides insight into its computational demands and efficiency. For an understanding of the application-specific energy consumption characteristics associated with each benchmark, we evaluated two divisions of inference benchmarks, datacenter and tiny.
Datacenter The blue bars in Figure 6 show the energy per inference sample and workload computation size for each MLPerf Inference benchmark on a large-scale datacenter system. ResNet [22, 55], a classic image classification model widely used as a baseline for computer vision (CV) tasks, consumes the least energy per inference at 8.7 mJ/Sample. As computer vision models get larger and more capable, the energy for a single inference also grows. RetinaNet [43] is an object detection model that uses multiple ResNet inferences for multiple classifications within a single image and thus a multiple order-of-magnitude growth in inference energy. The computer vision benchmark, 3D U-Net [7] is a new 3-dimensional medical image segmentation model that uses even more energy per inference due to the extra visual dimension and required precision.
Figure 6: Energy consumed per inference for each of the MLPerf Inference and MLPerf Tiny benchmarks. Each benchmark is also categorized by its workload type.
@@
We also evaluate two NLP benchmarks. RNN-T [31] is a speech recognition model designed to transcribe spoken language into text, while BERT [13] is a multipurpose language model capable of various tasks such as question answering. Despite their differences, these models have inference energy requirements similar to those of computer vision and recommendation models, likely due to their established architectures.
@@
It is important to note that autoregressive language models exhibit variability in their inference lengths. In MLPerf Inference v4.0, each of the two LLM benchmarks uses a different metric. GPTJ is reported in samples/second, while Llama2 is given in tokens/second. From version 4.1 onward, all LLM benchmarks will be reported in tokens/second, as is common in industry. However, for better alignment with how humans interact with LLMs and comparability with other MLPerf Inference benchmarks, we use samples/second to evaluate LLM performance for energy efficiency calculations in Fig. 6. To do this, we convert tokens into samples using an estimated median sequence length of 69 tokens/sample for GPTJ and 292 tokens/sample for Llama2, with an error margin due to sequence length variation, user pairs, and other factors.

### v2

@@ Abstract
Rapid adoption of machine learning (ML) technologies has led to a surge in power consumption across diverse systems, from tiny IoT devices to massive datacenter clusters. Benchmarking the energy efficiency of these systems is crucial for optimization, but presents novel challenges due to the variety of hardware platforms, workload characteristics, and system-level interactions. This paper introduces MLPerf® Power, a comprehensive benchmarking methodology with capabilities to evaluate the energy efficiency of ML systems at power levels ranging from microwatts to megawatts. Developed by a consortium of industry professionals from more than 20 organizations, coupled with insights from academia, MLPerf Power establishes rules and best practices to ensure comparability across diverse architectures. We use representative workloads from the MLPerf benchmark suite to collect 1,841 reproducible measurements from 60 systems across the entire range of ML deployment scales. Our analysis reveals trade-offs between performance, complexity, and energy efficiency across this wide range of systems, providing actionable insights for designing optimized ML solutions from the smallest edge devices to the largest cloud infrastructures. This work emphasizes the importance of energy efficiency as a key metric in the evaluation and comparison of the ML system, laying the foundation for future research in this critical area. We discuss the implications for developing sustainable AI solutions and standardizing energy efficiency benchmarking for ML systems.
@@ I Introduction
The advancement in ML capabilities has coincided with a notable increase in concerns about the power consumption of ML systems [76, 11, 75, 78, 65]. As the ecological impact and operational expenses of these systems continue to increase, they have become a significant concern for both industry and academia [82, 72, 80, 15]. The critical nature of this challenge highlights the need for a standardized and thorough benchmarking approach to measure the energy efficiency of ML systems, from small IoT devices to large-scale cloud infrastructures. This method must also guarantee accurate reproducibility, facilitate meaningful comparisons, and encourage advances in energy-efficient ML technologies.
@@
These diverse requirements emphasize the need for a comprehensive and adaptable power measurement methodology. To address the multifaceted challenges of measuring and comparing the energy efficiency of ML systems, we developed MLPerf Power, a comprehensive benchmarking methodology. This initiative is the result of collaborative efforts among 20 organizations actively participating in the MLPerf Training [46] and Inference [70] benchmarks, which have long served as the industry standard for measuring ML system performance. MLPerf Power extends this framework, recognizing that performance metrics alone are insufficient to enable the development of sustainable AI systems for the future.
@@ II Background, Motivation and Prior Work
This section establishes the requirements for benchmarking and measuring the energy efficiency of ML systems and demonstrates how our approach advances the field. We outline the key requirements and then gives an overview of the limitations of existing methods using Table I. MLPerf Power was specifically designed to meet all the identified requirements.
@@ III ML System Power Measurement Challenges
The goal of MLPerf is to be a system level performance benchmark for machine learning (ML) systems. Therefore, our goal while developing the power measurement methodologies for MLPerf was to develop a full system power measurement methodology that is applicable across the diverse hardware platforms that MLPerf supports and scalable across different MLPerf categories that target different system scales. This section highlights the challenges encountered while developing the MLPerf power measurement methodology.
@@ III-A Scale of ML Systems
Figure 2: The power consumption range across MLPerf divisions, highlighting the need for scalable power measurement.
@@
MLPerf supports a large diversity of submitters, each with their own telemetry and monitoring setup. However, measurement of interconnect power is challenging, as switches may not have as much telemetry to report power and cloud service providers often do not want to reveal their network topology to competitors. To address this, we need to develop a software-based approach that leverages estimation methods and existing telemetry systems, allowing for comprehensive power measurement without compromising security or requiring extensive hardware modifications.
Measuring cooling power fairly is another challenge. Air-cooled servers have fans whose power is measured when measuring node power, but immersion or direct liquid cooled systems pay their cooling cost at the datacenter level, making it difficult to attribute what fraction was used by the ML job under test. Measuring power consumption from cooling remains future work.
@@
Mobile MLPerf Mobile benchmarks the performance of mobile ML systems [32]. However, we currently do not support the measurement of power of mobile systems in MLPerf Power due to the complexity and variability of smartphones. Mobile devices comprise a wide range of components, including communication interfaces, sensors, and peripherals, each with significant fluctuations in power consumption patterns and active components[60]. Additionally, the battery-powered nature of these devices means they cannot be simply connected to an external power source for measurement without altering normal operation. Thermal throttling further complicates the relationship between power consumption and performance metrics [66]. We have piloted efforts using DC-based low-power analyzers, but we have not reached the point of standardization. Given these complexities, establishing a fair and transparent benchmark for measuring power consumption in battery-powered devices is a future goal for MLPerf Power.
@@ IV MLPerf Power Measurement Methodology
The methodology is designed to provide a comprehensive and standardized approach to measuring power consumption in ML systems across a wide range of scales, from microwatts to megawatts. This section outlines the core concepts, metrics, and philosophy behind the MLPerf power measurement approach.
@@ IV-A Common Principles Across All Segments
First, a fundamental tenet of MLPerf Power is the emphasis on measuring full system power consumption. This approach recognizes that ML workloads involve various components beyond the primary computing units. In datacenter systems, we account for compute nodes and interconnect fabric. As of today, we do not yet a methodology for including the power for the cooling devices and the storage nodes, but any other drives within the compute blades are accounted for. For edge devices, we consider the entire system-on-chip, including any peripherals that cannot be powered down during operation. In tiny systems, we measure the power of the entire device, including any always-on components. Figure 3 illustrates our measurement for the different scales of systems. Our measurements closely reflect the true energy cost of running ML workloads in real-world scenarios, rather than focusing solely on idealized component-level efficiencies.
Second, ML workloads typically consist of multiple phases: initialization, execution, and result assimilation. The MLPerf Power methodology emphasizes the precision of aligning power measurements with these phases, particularly focusing on the execution phase. This alignment is crucial because it allows for accurate attribution of power consumption to specific parts of the workload, ensures comparability across different systems, and provides insight into the energy efficiency of the core ML computations. To achieve this in tiny systems, we use hardware pins to signal the start and end of inference. In larger systems, we use software-based timestamps to demarcate the execution phase. For training workloads, we parse performance logs to determine the start and stop times of the run.
Third, unlike evaluating overall system performance, the execution phase is more important than end-to-end runtime in power measurement considerations. The execution phase represents the core ML computations, which are typically the most energy-intensive part of the workload. It also provides the most relevant power measurement data to draw actionable insights to improve the energy efficiency of ML systems where it matters most. For inference tasks we measure the power during the actual inference computations. For training tasks, we focus on the power consumed during the training iterations alone. Any of the preceding phases including data assimilation, data setup, offline and online pre-processing steps, are not accounted for. This approach, combined with full system power measurement (as explained in our first principle) and precise alignment with workload phases, forms the foundation of MLPerf Power’s methodology across all segments, from the smallest IoT devices to the largest datacenter installations.
@@
Finally, we employ a unified approach to evaluating energy efficiency on different scales and types of systems. For throughput benchmarks, such as datacenter and offline edge inference [70], we measure throughput in Samples/s and power consumption in Watts. We calculate energy efficiency as (Samples/s)/(Watts), yielding Samples/Joule. For latency benchmarks such as tiny inference [5], we use the inverse of energy in 1/Joules. We treat these two metrics as comparable, despite their different units, because the benchmarks fix different execution factors and thus require different performance evaluations. This standardized approach allows for meaningful comparisons across the diverse range of systems.
@@ IV-B MLPerf Inference Setup and Operation
The energy measurement setup for systems in the Tiny division is illustrated in Figure 4(a). To accurately measure such low-power devices, we utilize an I/O Manager that sits between the host and the System Under Test (SUT). The I/O Manager is implemented with an Arduino UNO board that captures timing information with resolution that the host (a Windows, Linux, or Mac) cannot. The I/O Manager, along with the level shifter, also electrically isolates the SUT from the host to avoid parasitic powering from the host’s signal lines, skewing the measurements. The I/O Manager provides a USB connection to the host and connects to the SUT’s Universal Asynchronous Receiver-Transmitter (UART) through an isolating level shifter, ensuring minimal power transfer.
To measure the energy, the host issues an infer command, which is relayed to the SUT via the I/O manager and includes a number of inferences to be performed. The SUT signals the beginning and end of the inference activity by toggling the Timestamp pin. This timing information is captured by the external energy monitor along with the current and voltage waveforms and transmitted to the host. The host then integrates the power over the indicated interval and divides by the number of inferences to calculate the energy per inference.
@@
Currently, the SUT collects performance log files that timestamp different phases of execution. To help us improve measurement accuracy, we employ a range mode that involves an initial run to determine the maximum current and voltage levels for a particular workload. Subsequent runs then use fixed analyzer ranges based on these observed peaks, allowing for higher accuracy within that range. For multi-node inference configurations, we adapt our methodology to handle multiple SUTs connected to a single analyzer or multiple analyzers connected to a single SUT, as shown in Figure 4(b). This flexibility allows us to accurately measure power consumption in more complex distributed inference scenarios.
@@
(a) Tiny System
(b) Multi-SUT Inference System
(c) Training System
Figure 4: Measurement diagrams for Tiny, Multi-SUT Inference, and Training systems.
@@ IV-C MLPerf Training and HPC Setup and Operation
Power measurement at the node level is the foundation of our approach, as shown in Figure 4(c). For each participating compute node, power measurements are taken using the submitter’s own telemetry systems, such as IPMI [48] or RedFish [21]. We recommend out-of-band measurement techniques to minimize interference with the ML job itself. In cases where individual node measurement is not feasible, power can be measured at the Power Distribution Unit (PDU) level, provided that all nodes supplied by that PDU are involved in the ML job. Submitters are required to thoroughly document the accuracy of their telemetry systems for review.
@@ IV-D Validation and Reporting
Quality Accuracy requirements form the foundation of our validation process. Although we recognize that the diverse range of systems and scales in ML workloads requires different approaches to power measurement, all must meet stringent accuracy standards as described by MLPerf Inference  [70] and Training [46]. For datacenter and high-performance computing systems, we require documentation of the telemetry systems used, including their accuracy specifications and calibration procedures. Edge systems using external power analyzers must employ SPEC-approved devices, known for their high accuracy and reliability. For TinyML, we mandate specialized micro-power instruments with target accuracy ratings [5].
@@
The HPC and training reporting and submission process is designed to more comprehensive, flexible, and contextually rich. Submitters are required to report power consumption data along with additional information that could affect these measurements, including details on cooling solutions, power management techniques, and environmental conditions during the testing. The process allows for reporting at either the PDU or node level, accommodating different measurement setups. Submitters run the benchmark with their chosen telemetry option, estimate switch power if necessary, and post-process power logs to conform to the MLPerf Logging Library format. For systems where certain components’ power consumption must be estimated, such as interconnect power in large-scale training sets, dedicated sections are provided in the reporting format to document the estimation methodology. The submission package must include the processed data in a specified directory structure, along with documentation on telemetry accuracy. This comprehensive approach ensures transparency and comparability between different submissions, while allowing for the flexibility needed to accommodate various system configurations and measurement approaches.
Review Our review process [52, 57] is rigorous and multilayered, designed to catch any inconsistencies or anomalies in the submissions. Upon receiving a submission, our review committee, composed of experts in ML systems and power measurement, examines the reported data and documentation. They verify that all required information is present, that the reported measurements comply with the accuracy requirements, and that the power consumption figures are consistent with the system specifications and workload characteristics.
Since the ecosystem is still evolving, in cases where novel measurement techniques or system architectures are involved, the review process may include additional steps. This might involve consultations with domain experts or requests for more detailed explanations of the measurement methodology. The goal is to ensure that innovative approaches can be included in the benchmarks while maintaining the high standards of accuracy and comparability that MLPerf is known for.
Transparency is a key principle of our validation and reporting process. All accepted submissions are made publicly available, along with their full documentation. This allows community scrutiny and fosters trust in the benchmark results. It also provides valuable data for researchers and engineers working to improve the energy efficiency of ML systems [19].
@@ V Results/Evaluation
We analyze this diverse range of data to gain an understanding of comparative trends across different workloads and system scales, shedding light on the areas of focus within the industry. MLPerf benchmarks span a diverse range of ML workloads, from resource-constrained keyword spotting to complex computer vision, sophisticated recommendation systems, and state-of-the-art large language models (LLMs) [47, 49]. With each version, we make changes to the MLPerf benchmarks to adapt to the changing ML landscape, such as updating datasets, adding new benchmarks, updating the network architectures, and refining our methodology and hyperparameters [69]. A notable example is the addition of LLMs in version 3.1 to address the growing interest in this area.
@@ V-C Workload-Specific Insights
Unlike raw power measurements, which are impacted by how submitters scale their hardware in relation to model size, energy consumption is agnostic to constant power or latency requirements. For this reason, we use the energy per benchmark inference to evaluate its computational demands and efficiency. We provide insights into the relationships between model size, workload characteristics, and energy consumption of each benchmark submission by evaluating two divisions of inference benchmarks, datacenter and tiny. Note that while there is no perfectly comparable metric for model size, we use total matrix accumulation (MAC) operations as an estimate. For an in-depth description of the workloads and deeper insight into the computation of each benchmark, refer to the MLPerf-Inference [70] and MLPerf-Tiny [5] papers.
Datacenter The blue bars in Figure 7 show the energy per inference sample and workload computation size for each MLPerf Inference benchmark on a large-scale datacenter system. ResNet [24, 59], a classic image classification model widely used as a baseline for computer vision (CV) tasks, consumes the least energy per inference at 8.7 mJ/Sample. As computer vision models get larger and more capable, the energy for a single inference also grows. RetinaNet [45] is an object detection model that uses multiple ResNet inferences for multiple classifications within a single image and thus a multiple order-of-magnitude growth in inference energy. The computer vision benchmark, 3D U-Net [7] is a 3-dimensional medical image segmentation model that uses even more energy per inference due to the extra visual dimension and required precision.
@@
We also evaluate two NLP benchmarks. RNN-T [33] is a speech recognition model designed to transcribe spoken language into text, while BERT [13] is a multipurpose encoder-only transformer language model capable of various tasks such as question answering. Despite their model architecture similarities, these workloads likely have much lower inference energy requirements than an LLM like GPT-J due to 1 order of magnitude fewer parameters and and 2 orders of magnitude fewer MAC computations.
Figure 7: Energy consumption and total MAC operations per inference for each MLPerf Inference and MLPerf Tiny benchmarks. Each benchmark is categorized by its workload type.
@@
It is important to note that autoregressive language models exhibit variability in their inference lengths. In MLPerf Inference v4.0, each of the two LLM benchmarks uses a different metric. GPT-J is reported in samples/second, while Llama2 is given in tokens/second. From version 4.1 onward, all LLM benchmarks will be reported in tokens/second, as is common in industry. However, for better alignment with how humans interact with LLMs and comparability with other MLPerf Inference benchmarks, we use samples/second to evaluate LLM performance for energy efficiency calculations in Fig. 7. To do this, we convert tokens into samples using an estimated median sequence length of 69 tokens/sample for GPT-J and 292 tokens/sample for Llama2, with an error margin due to sequence length variation, user pairs, and other factors.